OpenStudy (anonymous):

How would you factor this 2√12x4y3 ?

OpenStudy (hba):

\[2\sqrt{12x^4y^3},Right ?\]

OpenStudy (anonymous):

Yeah

OpenStudy (hba):

\[2\sqrt{2 \times 2 \times 3 \times x^2 x^2 \times y^2 \times y}\] Got this ?

OpenStudy (hba):

@keky123321 ?

OpenStudy (hba):

Please respond.

OpenStudy (anonymous):

he is sleeping i guess

OpenStudy (hba):

lol